OpenStudy (anonymous):

What are the solutions of 3x^2 + 15 = -6x?

OpenStudy (owlcoffee):

when we have: \[3x^2 +15=-6x\] We can transport it to the form: \[Ax^2 + Bx+C=0\] All we do is adding 6x to both sides: \[3x^2+6x+15=0\] we can divide the whole equality by 3 to get: \[x^2 + 2x +5=0\] adn now, we have to use the quadratic formula to find the solutions: \[x=\frac{ -2 \pm \sqrt{2^2 -4(1)(5)} }{ 2 }\]

OpenStudy (anonymous):

-1 ± i -1 ± 2i -2 ± i -2 ± 2i

OpenStudy (anonymous):

does are the answer choices

OpenStudy (owlcoffee):

okay, try operating the equality on my last posr. so you can find the value of x

OpenStudy (anonymous):

i dont know how to set this kinda problem up in all honesty i was just looking for the answer

OpenStudy (owlcoffee):

thing is, I am here to help you out, not to give you the answer, but I'll take another step for you: \[x=\frac{ -2\sqrt{2^2-4(1)(5)} }{ 2 }\] \[x=1 \pm \frac{ \sqrt{-16} }{ 2 }\] \[x= 1 \pm 2i\]

OpenStudy (owlcoffee):

sorry, it's: \[x=-1 \pm 2i\]

OpenStudy (anonymous):

well in this case i only needed the answer its the last problem of my test it wouldn't hurt to get one free answer when i have solved every last one of them but the problem i asked you but thanks anyway

OpenStudy (owlcoffee):

no problem.